Circular No. 6141

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CENTAURI 1995 S. Benetti, European Southern Observatory (ESO); M. Della Valle,

Dipartimento di Astronomia dell’Universitá di Padova; and S. Molendi, Istituto di Fisica Cosmica del CNR, Milan, report that inspection of a fully-reduced CCD spectrogram (range 500–900 nm, resolution 0.4 nm) obtained on Feb. 27.2 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) confirms this object to be a galactic nova caught during its early decline. The spectrum is dominated by strong emission lines of Hα, O I (multiplets 1, 4, 34), and Ca II (infrared triplet). Hα and O I (844.6 nm) have a FWZI of ∼ 6000 km/s and double-peaked profiles (peak-to-peak separations of ∼ 2000 km/s), with the redder component always stronger. Weaker emissions due to Fe II (multiplets 42, 46, 48, 49), Na I D, and He I (587.6 nm) are also detected.

Photometry by A. C. Gilmore, Mt. John University Observatory: Feb. 26.388 UT, V = 8.49, U − B = −0.67, B − V = +0.30, V − R = +1.24, V − I = +1.35 (airmass 1.67; comparison stars Cousins F218 and F205).

X-RAY NOVA 1993 IN OPHIUCHUS K. Borozdin, N. Alexandrovich, and R. Sunyaev, Space Research In-

stitute, Moscow, on behalf of the MIR-Kvant team, report: “This object (GRS 1716–249 = GRO J1719–24) continues to be bright according to observations made on Feb. 16 with the TTM instrument onboard MIR-Kvant. The flux from the source in the energy band 2–27 keV was equal to 275 ± 20 mCrab, which is the highest value detected by the TTM from this source since the main outburst in Sept. 1993 (IAUC 5878). The spectrum can be approximated by a power law with index 2.1 ± 0.3.”

SUPERNOVAE 1995A, 1995B A. V. Filippenko and A. J. Barth, University of California at Berkeley,

report that inspection of CCD spectra (range 420–700 nm, resolution 0.7 nm) obtained on Feb. 24 with the 3-m Shane reflector at the Lick Observatory reveals that SN 1995A (cf. IAUC 6131) is of type Ia, ∼ 2 months past maximum. The recession velocity, determined from strong emission lines in the parent galaxy, is 9400 km/s. Similar observations show that SN 1995B (cf. IAUC 6132) is of type II, with a recession velocity of 9900 km/s. The spectrum suggests that SN 1995B is still in its photospheric phase.

1995 February 27 Daniel W. E. Green